OpenStudy (anonymous):

A car traveling 24 m/s begins to decelerate at a constant rate of 5 m/s2. After how many seconds does the car come to a stop?

OpenStudy (agreene):

\[v(t) = 24ms^{-1}\] \[\alpha(t)=5ms^{-2}\] \[\int\limits_{}^{}\alpha(t)dt = v(t)\] \[\int\limits_{}^{} 5 dt=5t\] \[5t=24\] \[t=\frac{24}{5}s\] \[t=4.8s\]

OpenStudy (anonymous):

what webpage did you us e

OpenStudy (agreene):

what do you mean?